# The DECADE cosmic shear project III: validation of analysis pipeline using spatially inhomogeneous data

### v1

@@ 5.2 Robustness to nonlinear power spectrum and IA models
We have made certain choices in the modeling pipeline based on our best, current knowledge (which is informed by the latest cosmic shear analyses) on the validity/accuracy of each model. In particular, the nonlinear power spectrum and IA model are chosen based on arguments detailed in Section 3. In this section we check the shift in $S_{8}$ constraints if under alternative choices for these models. Note that in all cases, the changes are made only to the analysis pipeline; the simulated data vector is still generated using our fiducial setup.
Line 5 in Figure 3 shows the $S_{8}$ constraints when using HaloFit as our model for the nonlinear matter power spectrum. Using HaloFit to analyze a HMCode-generated data vector shifts the $S_{8}$ constraint to a slightly lower value, which is consistent with what is seen in e.g., Secco & Samuroff et al. (2022), and is because at fixed cosmology, the HaloFit model predicts slightly more power on nonlinear scales relative to HMCode (Mead et al. 2016, 2020). Moving onto IA model choices, line 7 in Figure 3 shows that the use of NLA (with two free parameters) tightens the cosmological constraints slightly ($\approx\!8\%$) relative to TATT, which has five free parameters. TATT is a more general model for intrinsic alignments and includes three additional parameters (see Table 1).
@@ 5.3 Impact of priors on nuisance parameters
We next check the impact of the different nuisance parameter marginalization on the final constraining power. Starting from the fiducial analysis setup, we either fix or broaden priors for the shear calibration $m$ and the mean redshift calibration $\Delta z$. We also run a variant where we fix the IA parameters. Line 9-13 of Figure 3 shows the different constraints from changing the priors on these nuisance parameters. The shear calibration uncertainties ($m$ priors) are subdominant, as narrowing/widening the parameter priors has negligible impact on the final constraints (line 9 and 10). The parameters associated with the redshift uncertainty show some small effect as fixing their values reduces the uncertainty in $S_{8}$ by $\approx\!8\%$ (line 11) and widening the prior by three times increases the uncertainty in $S_{8}$ by $\approx\!20\%$ (line 12). Finally, fixing all IA parameters improves the constraints ($\approx\!33\%$, line 13) relative to our fiducial case. These results indicate after our scale cuts are applied, our constraining power is not limited by systematic uncertainties from our shear and redshift calibrations, i.e., improving our precision on these systematic parameters will not significantly reduce the uncertainties on $S_{8}$. We are, however, limited by our uncertainties in the IA parameters. This is also explored in more detail in the data analysis of Paper IV.
@@ 5.4 Robustness to sampling choices and noise
As discussed in Section 4.3, all cosmology results for our tests (including those found below and those in Paper I) are obtained using the Nautilus (Lange 2023) sampler, whereas our main cosmology results will be extracted using the Polychord sampler (Handley et al. 2015) run on the high accuracy setting (Table 2). We show in line 1 and 14 of Figure 3 that under the fiducial settings, the results from Nautilus and Polychord are very consistent with each other both in the mean and the width of the posterior on $S_{8}$. Therefore, we use Nautilus to obtain the rest of the constraints presented in this work and all constraints in Paper IV. We also show in line 15 the same analysis run with the Multinest sampler, which show slightly underestimated error bars (by about 7% in $S_{8}$). This has been pointed out in previous work (Lemos & Weaverdyck et al. 2023) and we confirm that result here.
@@ 6 Impact of spatial inhomogeneities
As discussed previously, the DECADE cosmic shear catalog is constructed using image data from a wide variety of community-led campaigns, which in turn results in a variety of observing conditions and image quality. This is in contrast to other dedicated lensing surveys — such as DES, KiDS, and Subaru Hyper Suprime Cam (HSC, Miyazaki et al. 2018) — where homogeneity in image quality and depth is a target of the survey observing strategy.
We reiterate a subtlety discussed above in Section 4.2: the impact of survey inhomogeneity on source galaxy number counts, $n_{\rm gal}$, is largely alleviated through our sample selection (a cut of $m_{i}<23.5$), and so the remaining, predominant impact of inhomogeneous image quality is on the measured galaxy properties. Figure 1 illustrates this point. This does not mean the DECADE survey is “as homogeneous” as DES Y3. The observing conditions of the survey — which we will describe below in Section 6.2 — will impact the distribution of measured galaxy properties in a spatially dependent manner. This can still impact the measurements and therefore, the final cosmology constraints.
@@ 6.1 Methodology
Figure 4: The constraints on $S_{8}$ from different subsets of the DECADE footprint, while consistently re-deriving all calibrations, redshift distributions, covariance matrices etc. The first, “Fiducial” constraint is using data the entire dataset. The rest, from top to bottom, split the survey based on: airmass, seeing (PSF width), DCR effects in celestial/ellipticity coordinates, the east/west regions, the $\rm Dec.\gtrless-30$ region where photometric calibrations used Pan-STARRS/SkyMapper, the sky background brightness and variations, exposure time, magnitude limit, number of exposures, interstellar extinction coefficients, and Gaia stellar number density. Analyses with $X^{+}$ ($X^{-}$) use the area with higher (lower) values than the median. The data vectors in each subset have essentially independent noise realizations and so the cosmology constraints can exhibits relative shifts of up to $3\sigma$ due to just statistical fluctuations. We show the parameter constraints relative to the fiducial estimate, with thick and thin horizontal lines denoting the $68\%$ and $95\%$ intervals for each constraint, respectively. For visibility reasons, we do not show the $95\%$ interval for the “SkyMapper” subset alone. Most (all) deviations are within $1\sigma$ ($2\sigma$) of the fiducial results. The constraints from a subset and its complement are always within $2\sigma$ as well. No split exhibits shifts in cosmology that are larger than the expected variation from shape noise and cosmic variance. The footprints corresponding to the splits are shown in Figure 9.
Our end-to-end test involves splitting the source-galaxy sample into subsets (usually halves) based on some criteria (see Section 6.2 and 6.3) and redoing all measurement and calibration steps; this includes the shear calibration (Paper I), redshift estimation/calibration (Paper II), analytical covariance generation, data vector measurements, and cosmological inference (this work). We consider splits both on the survey area and on the galaxy properties. The former is a clean split on the inhomogeneity in the survey image quality, whereas the latter can isolate subsets of the catalog that may be more susceptible to systematics of a specific kind. We briefly detail all the steps in the pipeline that are re-run for each subset:
@@
Shear Calibration. As described above in Section 3.3, all cosmology analyses of weak lensing require some calibration of the multiplicative bias, $m$. This is often done using image simulations of the survey, which is also the approach in DECADE as detailed in Paper I. We use the existing, simulated catalog and repeat the estimates of $m$, while incorporating the new selection function (which defines the split) that is being applied to the data.
Redshift estimation. For our cosmology analysis, we estimate the $n(z)$ using Self-Organizing Maps (SOMs; Kohonen 1982, 2001; Carrasco Kind & Brunner 2014; Masters et al. 2015) and the technique is denoted as SOMPZ (Buchs et al. 2019; Myles & Alarcon et al. 2021). One aspect of this method is classifying the source galaxies into a set of phenotypes, $\hat{c}$ (see Paper II). For these split tests, we do not reassign galaxies to a given phenotype and simply use the existing classification. [^6] This choice induces no bias in the derived $n(z)$ and only impacts their potential “optimality”; namely the overlap between the redshifts distributions of two tomographic bins. However, we redo the selection of galaxies into four tomographic bins, and the estimation of the $n(z)$ per tomographic bin. This includes re-estimating the transfer function — through the Balrog synthetic source catalog built for DECADE— to connect the true properties of a galaxy with its observed properties. In all cases, we apply the same area/property selections to the galaxy samples. We generate a large number of $n(z)$ realizations, as is done in Paper II, and characterize the uncertainty in the mean using the same method as described in that work. We do not repeat any measurements of the clustering redshifts; these measurements are only used a cross-check in Paper II and do not inform the fiducial $n(z)$ used in our analysis.
Covariance matrix. We estimate new $n_{\rm eff}$ and $\sigma_{e}$ quantities (effective number density and shape noise of source galaxies) for the sample in our subset. We also recompute the geometry of the survey footprint, incorporating our new selections, and use it in conjunction with the new $n(z)$ to generate our analytic covariance matrix. This is done through CosmoCov, as is described above in Section 4.2.
Data vector and cosmology. We re-measure the data vector using only galaxies in the chosen part of the footprint, while using the new tomographic bin separations and the new shear weights. During our internal tests, the data vectors from this step were all blinded using the same random seed used on the data (Paper IV). However, our results presented below are the constraints re-evaluated after unblinding the data vector.
@@
number of $\text{number of }\sigma=\frac{[S_{8}]_{1}-[S_{8}]_{2}}{\sqrt{\sigma([S_{8}]_{1})^{2}+\sigma([S_{8}]_{2})^{2}}},$ (11)
where $[S_{8}]_{i}$ and $\sigma([S_{8}]_{i})$ are the mean and standard deviation of the posterior for constraint $i$. Unlike the previous tests, which involved consistency between analysis choices operating on the exact same data vector, this test uses data vectors with different noise realizations that are essentially uncorrelated, [^7] The shape-noise realizations are completely uncorrelated between two subsets that use different galaxies in their measurements. However, the cosmic shear measurements can still be correlated due to large-scale density fluctuations that are shared by subsets spanning adjacent regions of the sky. This is the “super sample covariance” discussed in Section 4.2. We note that shape noise is the dominant contributor to the measurement covariance. since they access different subsets of galaxies from different subregions of the DECADE footprint. For this reason, it is possible for cosmology constraints from a given split to be different purely due to random fluctuations (where the fluctuations are dominated by shape noise, but will also have some cosmic variance component). Thus, our null test is satisfied if the $S_{8}$ constraints from a given subset are within $3\sigma$ of the constraints from the complementary subset. Though, in practice, we find all our consistency estimates are comfortably below this threshold. Here, we define complementary subset as “all galaxies not in the subset”.
@@ 6.2 Survey property splits & results
Figure 5: The distribution of magnitude limits in the $i$-band (from a HealPix map of $\texttt{NSIDE}=4096$) for the two subsets of the DECADE footprint split by magnitude limit, and for the DES Y3 footprint. The Maglim+ region has twice the width of the other two samples. The ranges in the figure denote the $1\%$ and $99\%$ values of the distribution. There is a slight overlap in the two DECADE distributions as we plot values from the $\texttt{NSIDE}=4096$ maps, whereas the selection mask is defined using the downgraded $\texttt{NSIDE}=1024$ maps (see Section 6.2 for details).
@@
Figure 6: Similar to Figure 4 but now splitting the shear catalog on a variety of galaxy properties. From top to bottom, we split the catalog based on: $r-z$ color, magnitude in $riz$ bands, signal-to-noise, the object size (in arcsec2), the PSF size (in arcsec2), and the ratio of object size to PSF size. All quantities are taken from the Metacalibration catalog.
@@
We first split on a number of different area-dependent quantities, most of which are informed by the survey observing conditions. In Paper I (see their Figure 13), we check that the measured galaxy shapes are not correlated with a wide variety of survey properties: the seeing (PSF width), airmass, differential chromatic refraction (DCR) in $e_{1}$ and $e_{2}$ directions, DCR in the right ascension and declination directions, the magnitude limit, exposure time, number of exposures, the standard deviation of the sky background, and the sky brightness. We use this same set of maps in our test here, and also add a number of additional splits defined below. Every survey property map has four variants, for the $griz$ bands. The $g$-band data is not used in any of our lensing analysis. Of the remaining three, we choose the $i$-band maps for all our inhomogeneity tests.
[…]
@@
Figure 4 shows the results from all fifteen criteria, which corresponds to thirty different subsets of our footprint. The splits across survey properties often fill our full footprint in a somewhat patchy manner (see Figure 9), whereas the splits on geometry are distinct, internally contiguous footprints. In all cases, the data in the subsets will have independent realizations of the shape noise. The results in Figure 4 show that for any chosen subset, the constraints from the subset and its complement are within $2\sigma$ of each other. The blue set of points, which are all PSF-related, are correlated with each other, as are the purple set of points, which are all related to image depth. In summary, we find that all thirty splits are within $2\sigma$ of the fiducial result, and each subset is within $2\sigma$ of its complement, which passes our $3\sigma$ criteria.
@@
Area-dependent $n(z)$: While the above tests do not explicitly probe the impact of spatially correlated variations on the $n(z)$ estimates, they still implicitly check this effect. In particular, the regions of the sky with a higher magnitude limit are also regions with a significantly larger variation in the magnitude limit. Figure 5 shows distributions of these limits, and lists the 1% and 99% values of the distributions. The Maglim+ distribution is nearly twice as a wide as the Maglim- and DES Y3 distributions. Thus, any spatially varying $n(z)$ effects will be even more prominent in the Maglim+ region. The consistency of the cosmology from the Maglim+ and Maglim- regions (the constraints are within $1.2\sigma$ as shown in Figure 4) suggests our final cosmology result is not significantly impacted by such area-dependent redshift effects.

### v2

@@ 5.2 Robustness to nonlinear power spectrum and IA models
We have made certain choices in the modeling pipeline based on our best, current knowledge (which is informed by the latest cosmic shear analyses, e.g., Secco & Samuroff et al. (2022); Amon et al. (2022); Asgari et al. (2021); DES and KiDS Collaborations et al. (2023); Li et al. (2023)) on the validity and expected accuracy of each model. In particular, the nonlinear power spectrum and the IA model are chosen based on arguments detailed in Section 3. In this section we check the shift in $S_{8}$ constraints if under alternative choices for these models. Note that in all cases, the changes are made only to the analysis pipeline; the simulated data vector is still fixed to our fiducial setup.
Line 5 in Figure 3 shows the $S_{8}$ constraints when using HaloFit as our model for the nonlinear matter power spectrum. Using HaloFit to analyze a HMCode-generated data vector shifts the $S_{8}$ constraint to a slightly lower value, which is consistent with what is seen in e.g., Secco & Samuroff et al. (2022). The shift is because at fixed cosmology, the HaloFit model predicts slightly more power on nonlinear scales relative to HMCode (Mead et al. 2016, 2020). For the IA model comparison, line 7 in Figure 3 shows that the use of NLA (with two free parameters) tightens the cosmological constraints slightly ($\approx\!8\%$) relative to TATT, which has five free parameters. TATT is a more general model for intrinsic alignments and includes three additional parameters (see Table 1). The reduced flexibility of the NLA model (relative to that of TATT) restricts the parameter space and tightens the cosmology constraints.
@@ 5.3 Impact of priors on nuisance parameters
We next check the impact of the different nuisance parameter marginalization on the final constraining power. Starting from the fiducial analysis setup, we either fix or broaden priors for the shear calibration $m$ and the mean redshift calibration $\Delta z$, relative to the fiducial setup (Table 1). We also run a variant where we fix the IA parameters. Lines 9-13 of Figure 3 shows the different constraints from changing the priors on these nuisance parameters. The shear calibration uncertainties ($m$ priors) are a subdominant contribution to the final cosmology constraints, as narrowing/widening these parameter priors has negligible impact on the final constraining power (line 9 and 10). The parameters associated with the redshift uncertainty show some small effect as fixing their values reduces the uncertainty in $S_{8}$ by $\approx\!8\%$ (line 11) and widening the prior by three times increases the uncertainty in $S_{8}$ by $\approx\!20\%$ (line 12). [^8] Figure 3 shows that broadening the redshift calibration uncertainties, also shifts the mean of the $S_{8}$ marginal posterior away from the input value by around $0.5\sigma$ of the posterior width. This is a manifestation of projection effects, which can occur as the prior volume is expanded. We note that such projection effects are not present in our fiducial analysis (Line 1), where the redshift calibration uncertainties are not artificially broadened. Finally, fixing all IA parameters (Line 13) improves the constraints by 23% (13%) relative to our fiducial (NLA) case. These results indicate that, after applying scale cuts, the primary limitation on the constraining power of cosmology parameters is the uncertainties in the intrinsic alignment model — and not from the shear and redshift calibrations. The latter limitation from IA uncertainties is also explored in more detail in the data analysis of Paper IV.
@@ 5.4 Robustness to sampling choices and noise
As discussed in Section 4.3, all cosmology results for our tests (including those found below and those in Paper I) are obtained using the Nautilus (Lange 2023) sampler, whereas our main cosmology results will be extracted using the Polychord sampler (Handley et al. 2015) run on the high accuracy setting (Table 2). We show in line 1 and 14 of Figure 3 that under the fiducial settings, the results from Nautilus and Polychord are very consistent with each other both in the mean and the width of the posterior on $S_{8}$. The means are shifted by less than $0.05\sigma$ of each other and the posterior widths are within $4\%$ of each other. We also show in line 15 the same analysis run with the Multinest sampler, which show slightly underestimated error bars (by about 7% in $S_{8}$). This has been pointed out in previous work (Lemos & Weaverdyck et al. 2023) and we confirm that result here.
In summary, the Nautilus sampler is consistent with constraints from Polychord, where the latter is a widely used sampler in DES analyses and other works. We therefore adopt Nautilus for all remaining constraints presented in this work and for those presented in Paper IV.
@@ 6 Impact of spatial inhomogeneities
We now quantify the impact of any spatially varying effects on our cosmology constraints, using a fully data-driven approach. As discussed previously, the DECADE cosmic shear catalog is constructed using image data from a wide variety of community-led campaigns, which in turn results in a variety of observing conditions and image quality. This is in contrast to other dedicated lensing surveys — such as DES, KiDS, and Subaru Hyper Suprime Cam (HSC, Miyazaki et al. 2018) — where homogeneity in image quality and depth is a target of the survey observing strategy (e.g., Neilsen et al. 2019; Kuijken et al. 2019; Aihara et al. 2022).
We reiterate a subtlety discussed above in Section 4.2: the impact of survey inhomogeneity on source galaxy number counts, $n_{\rm gal}$, is largely alleviated through our sample selection (a cut on apparent magnitude $m_{i}<23.5$). Figure 1, and its associated discussion in Section 4.2, illustrates this point. However, this does not mean the DECADE survey is “as homogeneous” as DES Y3. The observing conditions of the survey — which we will describe below in Section 6.2 — will still impact the distribution of measured galaxy properties in a spatially dependent manner. For example, regions of the sky with deeper imaging (higher magnitude depth) will contain galaxies whose properties are measured to better precision. Such an effect can still impact the measurements of shear, redshifts, etc., and therefore, the final cosmology constraints.
@@ 6.1 Methodology
Figure 4: The constraints on $S_{8}$ from different subsets of the DECADE footprint, while consistently rederiving all calibrations, redshift distributions, covariance matrices, etc. The first, “Fiducial” constraint is using data the entire dataset. The rest, from top to bottom, split the survey based on: airmass, seeing (PSF width), differential chromatic refraction (DCR) effects in the R.A., Dec., and $e_{1,2}$ coordinates, the east/west regions, the $\rm Dec.\gtrless-30$ region where photometric calibrations used Pan-STARRS/SkyMapper, the sky background brightness and variations, exposure time, magnitude limit, number of exposures, interstellar extinction coefficients, and Gaia stellar number density. Analyses with $X^{+}$ ($X^{-}$) use the area with higher (lower) values than the median. The data vectors in each subset have essentially independent noise realizations and so the cosmology constraints can exhibit relative shifts of up to $3\sigma$ due to just statistical fluctuations. We show the parameter constraints relative to the fiducial estimate, with thick and thin horizontal lines denoting the $68\%$ and $95\%$ intervals for each constraint, respectively. For visibility reasons, we do not show the $95\%$ interval for the “SkyMapper” subset alone. Most (all) deviations are within $1\sigma$ ($2\sigma$) of the fiducial results. The constraints from a subset and its complement are always within $2\sigma$ as well. No split exhibits shifts in cosmology that are larger than the expected variation from shape noise and cosmic variance. The footprints corresponding to the splits are shown in Figure 9.
Our end-to-end test involves splitting the source-galaxy sample into subsets (usually halves) based on some criteria (see Section 6.2 and 6.3) and redoing all measurement and calibration steps; this includes the shear calibration (Paper I), redshift estimation/calibration (Paper II), analytical covariance generation, data vector measurements, and cosmological inference (this work). We consider splits both on the survey area and on the galaxy properties. The former is a spatial split defined using the inhomogeneity in the survey image quality, whereas the latter can isolate subsets of the catalog that may be more susceptible to systematics of a specific kind. We briefly detail all the steps in the pipeline that are rerun for each subset:
@@
Shear Calibration. As described above in Section 3.3, all cosmology analyses of weak lensing require some calibration of the multiplicative bias $m$. This is often done using image simulations of the survey, which is also the approach in DECADE as detailed in Paper I. We use the existing, simulated catalog and repeat the estimates of $m$, while incorporating the new selection function (which defines the split) that is being applied to the data.
Redshift estimation. For our cosmology analysis, we estimate the $n(z)$ using self-organizing maps (SOMs; Kohonen 1982, 2001; Carrasco Kind & Brunner 2014; Masters et al. 2015) and the technique is denoted as SOMPZ (Buchs et al. 2019; Myles & Alarcon et al. 2021). One aspect of this method is classifying the source galaxies into a set of phenotypes, $\hat{c}$ (see Paper II). For these split tests, we do not reassign galaxies to a given phenotype and simply use the existing classification. [^9] This choice induces no bias in the derived $n(z)$ and only degrades their potential “optimality”; namely the overlap between the redshift distributions of two tomographic bins. We do not quantify this effect in our work. However, we redo the selection of galaxies into four tomographic bins, and the estimation of the $n(z)$ per tomographic bin. This includes re-estimating the transfer function — through the Balrog synthetic source catalog built for DECADE (Paper II) — to connect the true properties of a galaxy with its observed properties. In all cases, we apply the same area/property selections to the galaxy samples. We generate a large number of $n(z)$ realizations, as is done in Paper II, and characterize the uncertainty in the mean using the same method as described in that work. We do not repeat any measurements of the clustering redshifts; these measurements are only used a cross-check in Paper II and do not inform the fiducial $n(z)$ used in our analysis.
Covariance matrix. We estimate new $n_{\rm eff}$ and $\sigma_{e}$ quantities (effective number density and shape noise of source galaxies, respectively) for the sample in our subset. We also recompute the geometry of the survey footprint, incorporating our new selections, and use it in conjunction with the new $n(z)$ to generate our analytic covariance matrix. This is done through CosmoCov, as is described above in Section 4.2.
Data vector and cosmology. We remeasure the data vector using only galaxies in the chosen part of the footprint, while using the new tomographic bin assignments and the new shear weights. During our internal tests, the data vectors from this step were all blinded using the same random seed used on the data (Paper IV). However, our results presented below are the constraints re-evaluated after unblinding the data vector.
@@
$\text{number of }\sigma=\frac{[S_{8}]_{1}-[S_{8}]_{2}}{\sqrt{\sigma([S_{8}]_{1})^{2}+\sigma([S_{8}]_{2})^{2}}},$ (13)
where $S_{8}]_{i}$ and $\sigma([S_{8}]_{i})$ are the mean and standard deviation of the posterior for constraint $i$, respectively. Unlike the previous tests, which involved consistency between analysis choices operating on the exact same data vector, this test uses data vectors with different noise realizations that are essentially uncorrelated, [^10] The shape-noise realizations are completely uncorrelated between two subsets that use different galaxies in their measurements. However, the cosmic shear measurements can still be correlated due to large-scale (cosmological) density fluctuations that are shared by subsets spanning adjacent regions of the sky. We note that shape noise is the dominant contributor to the measurement covariance. since they access different subsets of galaxies from different subregions of the DECADE footprint. For this reason, it is possible for cosmology constraints from a given split to be different purely due to random fluctuations (where the fluctuations are dominated by shape noise, but will also have some cosmic variance component). Thus, our null test is satisfied if the $S_{8}$ constraints from a given subset are within $3\sigma$ of the constraints from the complementary subset. Though, in practice, we find all our consistency estimates are comfortably below this threshold. Here, we define complementary subset as “all galaxies not in the subset”. In practice, there can be mild correlations between a subset and its complement. We have quantified these in two limiting cases and verified the standard score estimate, using Equation (13), is changed by 10-15% at most. See Appendix B for more details.
@@ 6.2 Survey property splits & results
Figure 5: The distribution of magnitude limits in the $i$-band (from a HealPix map of $\texttt{NSIDE}=4096$) for the two subsets of the DECADE footprint split by magnitude limit, and for the DES Y3 footprint. The ranges in the figure denote the $1\%$ and $99\%$ values of the distribution. The Maglim+ region (blue) has twice the width — defined as the difference between the 99% and 1% intervals shown in the Figure — compared to those of the other two samples. There is a slight, minimal overlap in the two DECADE distributions as we plot values from the $\texttt{NSIDE}=4096$ maps, whereas the selection mask is defined using the downgraded $\texttt{NSIDE}=1024$ maps (see Section 6.2 for details). The overlap has no impact on our qualitative and quantitative discussions.
@@
Figure 6: Similar to Figure 4 but now splitting the shear catalog on a variety of galaxy properties. From top to bottom, we split the catalog based on: rzr-z color, magnitude in rizriz bands, signal-to-noise, the object size (in arcsec2), the PSF size (in arcsec2), and the ratio of object size to PSF size. All quantities are taken from the Metacalibration catalog.
@@
We first split the sample on a number of different area-dependent quantities, most of which are informed by the survey observing conditions. In Paper I (see their Figure 13), we check that the measured galaxy shapes are not correlated with a wide variety of survey properties: the seeing (PSF width), airmass, differential chromatic refraction (DCR) in $e_{1}$ and $e_{2}$ directions, DCR in the right ascension (R.A.) and declination (Dec.) directions, the magnitude limit, exposure time, number of exposures, the standard deviation of the sky background, and the sky brightness. We use this same set of maps in our test here, and also add a number of additional splits defined below. Every survey property map has four variants, one for each of the grizgriz photometric bands. The $g$-band data is not used in any of our lensing analysis (see Paper I for more details). Of the remaining three bands, we choose the $i$-band maps for all our inhomogeneity tests, but note than the $r$ and $z$ bands would be equally viable choices.
We also split the footprint based on the Gaia stellar density (Gaia Collaboration et al. 2021) and interstellar extinction maps from Schlegel et al. (1998), which were used in Paper I to define our foreground mask (see discussion in Section 2.1). In all these cases, we split the footprint by sampling the map value for every object in our source-galaxy sample, finding the median value for the sample, and splitting the area of the sky based on that median. This ensures the splits contain an equal number of galaxies. The maps are downsampled to $\texttt{NSIDE}=4096\rightarrow 1024$ to improve the overall contiguousness of the selection mask by reducing the presence of disconnected patches in the survey mask. The remaining two split definitions are purely geometric (selections on $\rm R.A.$ and $\rm Dec.$ alone). First, we split our survey on ${\rm Dec.}\gtrless-30$ as all CCD images in regions north (south) of this line used the Pan-STARRS-1 (SkyMapper) measurements — as found in the ATLAS RefCat2 reference catalog (Tonry et al. 2018) — for their photometric calibrations (see Tan et al. 2024, for more details). Second, we split the footprint into an eastern and western footprint, split on ${\rm R.A.}\gtrless 180$. [^11] We assign the entire SkyMapper-calibrated region (${\rm DEC.}<-30$) to the western footprint as both regions were part of the same, internal image processing run. Altogether, we define fifteen different ways of splitting the footprint, with each relying on a different survey observing condition.
@@
Figure 4 shows the results from all fifteen criteria, which corresponds to thirty different subsets of our footprint. The splits across survey properties often fill our full footprint in a somewhat patchy manner (see Figure 9), whereas the splits on geometry are distinct, internally contiguous footprints. In all cases, the data in the subsets will have essentially independent realizations of the shape noise (Appendix B). The results in Figure 4 show that for any chosen subset, the constraints from the subset and its complement are within $2\sigma$ of each other. Here, we have used the metric defined in Equation (13). In summary, we find that all thirty splits are within $2\sigma$ of the fiducial result, and each subset is within $2\sigma$ of its complement, which passes our $3\sigma$ criteria.
@@
Area-dependent $\boldsymbol{n(z)}$: While the above tests do not explicitly probe the impact of spatially correlated variations on the $n(z)$ estimates, they still implicitly check this effect. In particular, the regions of the sky with a higher magnitude limit are also regions with a significantly larger variation in the magnitude limit. If the spatial variation of the magnitude limit has an impact on the true $n(z)$ of our source galaxy sample, then the spatial variation of the $n(z)$ will have a larger amplitude if the variation in the magnitude limit is increased. Figure 5 shows distributions of the magnitude limits under the two subsets, and lists the 1% and 99% values of the distributions. These values show that the Maglim+ distribution is nearly twice as a wide as the Maglim- and DES Y3 distributions. Thus, any spatially varying $n(z)$ effects will be even more prominent in the Maglim+ region. The consistency of the cosmology from the Maglim+ and Maglim- regions (the constraints are within $1.2\sigma$ as shown in Figure 4) suggests our final cosmology result is not significantly impacted by such area-dependent redshift effects.